# From shape to fate: making bacterial swarming expansion predictable

### v1

@@ 4.4 Morpher
Sequence model variants with matched capacity. The temporal module is instantiated using one of four architectures: a vanilla RNN [13], GRU [10], LSTM [16], or a Transformer encoder [42]. All variants share the same spatial encoder–decoder backbone and operate in the same latent space. Across architectures, temporal information is summarized into a single terminal latent state that serves as the sole conditioning variable for future prediction, enforcing a consistent past-to-state formulation. The recurrent variants use three recurrent layers, while the Transformer uses three stacked encoder blocks. No model introduces future tokens, decoder-side attention, or changes to the latent dimensionality.
Morphon memory integration. To regulate how past morphology influences future predictions, Morphon retrieves informative historical states from the latent history via cross-attention, using a learnable query derived from the current state. The retrieved representation is fused into the evolving temporal state through a learnable gate $\alpha\in(0,1)$, balancing historical memory against the current predictive trajectory.
Multi-scale decoding and feedback. A multi-scale decoder mirrors the encoder hierarchy to progressively restore spatial resolution while reinjecting intermediate features retained during encoding. This design preserves peripheral protrusions and fine-scale curvature patterns during reconstruction. During training, the predicted mask is re-encoded and fed back for subsequent steps, matching the inference-time autoregressive procedure.
Sequence structuring. Observation length $T_{\mathrm{obs}}$ and forecasting horizon are treated as tunable hyperparameters. Sequences are constructed by uniformly subsampling frames with a fixed stride and partitioned into observation and prediction segments under predefined observation–prediction ratios.
Loss. Supervision is applied at every predicted time step within the forecasting horizon. Let $T$ denote the number of predicted frames. The training objective is defined as a temporal average over the prediction window:
$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T}\sum_{t=1}^{T}\Big(\lambda_{\mathrm{focal}}\,\mathcal{L}_{\mathrm{focal}}^{(t)}+\lambda_{\mathrm{softIoU}}\,\mathcal{L}_{\mathrm{softIoU}}^{(t)}+\lambda_{\mathrm{boundary}}\,\mathcal{L}_{\mathrm{boundary}}^{(t)}\Big).$
Focal loss addresses foreground–background imbalance by up-weighting hard pixels, while soft IoU enforces region-level overlap. A boundary-aware loss further penalizes discrepancies along the colony front; it is computed using a GPU-efficient morphological gradient approximation and evaluated at one-quarter of the original spatial resolution. Loss terms are averaged uniformly over the prediction horizon, and the relative weights are fixed across experiments.
Interpretability as a dynamical model. By expressing each predicted morphology as a causal progression from prior front states, Morpher provides a data-driven dynamical representation of swarming expansion, enabling analysis of curvature perturbation propagation over time.
4.5 Evaluation metrics for colony front segmentation
We evaluated segmentation performance using four complementary metrics: mAP50:95, the Dice coefficient, the number of learnable parameters, and inference speed measured in frames per second (FPS). Among these, mAP50:95 serves as the primary metric, as it provides a comprehensive measure of segmentation quality across a range of localization tolerances. Following the COCO evaluation protocol, AP is first computed at multiple Intersection–over–Union (IoU) thresholds, and mAP50:95 is obtained by averaging across IoU thresholds ranging from 0.50 to 0.95, with a step size of 0.05:
$\mathrm{mAP}_{50:95}=\frac{1}{10}\sum_{\tau\in\{0.50,0.55,\dots,0.95\}}\left[\frac{1}{101}\sum_{i=0}^{100}P\!\left(\tfrac{i}{100};\,\tau\right)\right],$
where $P(r;\tau)$ denotes the interpolated precision at recall level $r$ for IoU threshold $\tau$.
In addition to mAP, segmentation accuracy is quantified using the Dice coefficient, which measures the overlap between predicted and ground-truth masks. For a predicted mask $\hat{Y}$ and a ground-truth mask $Y$, the Dice coefficient is defined as
$\mathrm{Dice}=\frac{2\,|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|},$
with values ranging from 0 to 1, where higher scores indicate better spatial agreement.
To assess model complexity, we report the total number of learnable parameters, which indicates the representational capacity and computational burden of the segmentation network. To evaluate practical efficiency, we additionally measure inference throughput in FPS, defined as the number of images the model can process per second under the same hardware configuration.
4.6 Evaluation metrics for morphological forecasting
Forecasting accuracy must be judged not only by per–frame agreement, but also by how faithfully the predicted colony advances and organizes its growth direction over time. Accordingly, our evaluation considers both the spatial fidelity of the predicted masks and boundaries, and the temporal consistency of the advancing front, including its radial expansion speed and the coherence of its directional variation around the colony rim. All symbols and operations used in the metrics below are defined explicitly.
To quantify mask fidelity over time, let $\widehat{Y}_{t}$ and $Y_{t}$ denote the predicted and true masks at time $t$, and let $\partial\widehat{Y}_{t}$ and $\partial Y_{t}$ denote their corresponding boundaries. We report the mean Intersection over Union (mIoU), defined as the temporal average of the IoU used in the AP protocol:
$\mathrm{mIoU}=\frac{1}{T}\sum_{t=1}^{T}\frac{\lvert\widehat{Y}_{t}\cap Y_{t}\rvert}{\lvert\widehat{Y}_{t}\cup Y_{t}\rvert}$.
To evaluate boundary placement in a manner sensitive to fine-scale morphological deviations, we use the symmetric Hausdorff distance,
$d_{\mathrm{H}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{\max_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|,\;\max_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\Big\},$
as well as its 95th–percentile variant,
$d_{\mathrm{H}_{95}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{P_{95}\!\big(\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|\big)_{y\in\partial Y_{t}},\;P_{95}\!\big(\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\big)_{\hat{y}\in\partial\widehat{Y}_{t}}\Big\},$
which downweights sparse, extreme outliers arising from transient high–curvature growth. Finally, the average symmetric surface distance provides a measure of mean bidirectional nearest–boundary discrepancy,
$d_{\mathrm{ASSD}}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\lvert\partial\widehat{Y}_{t}\rvert+\lvert\partial Y_{t}\rvert}\left(\sum_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|+\sum_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\right).$
To evaluate whether the predicted colony reproduces the correct front propagation dynamics, we measure the radial expansion speed along $K$ uniformly sampled angular directions $\{\theta_{k}\}_{k=1}^{K}$ from the colony centroid. In our experiments, we set $K=720$, which corresponds to a sampling interval of $0.5^{\circ}$, which provides sufficiently fine angular resolution to capture the subtle, high–curvature perturbations that arise during dense–finger instabilities. Let $r(\theta_{k},t)$ and $\widehat{r}(\theta_{k},t)$ denote the ground–truth and predicted radial distances at time $t$. The corresponding expansion velocities are
$v(\theta_{k},t)=\frac{r(\theta_{k},t)-r(\theta_{k},t-\Delta t)}{\Delta t},\qquad\widehat{v}(\theta_{k},t)=\frac{\widehat{r}(\theta_{k},t)-\widehat{r}(\theta_{k},t-\Delta t)}{\Delta t}.$
The overall accuracy of front advancement is quantified by the mean root–mean–square error,
$\mathrm{RMSE}=\frac{1}{T-1}\sum_{t=2}^{T}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\big(\widehat{v}(\theta_{k},t)-v(\theta_{k},t)\big)^{2}}.$
Beyond matching the average speed, we also assess whether the temporal fluctuation strength of the radial velocity field is preserved. To this end, we compute the Temporal Consistency Index (TCI) over sliding windows $w=1,\dots,W$ of fixed length $L=4$ radius frames, corresponding to three consecutive velocity steps. For each window and direction, we compare the temporal standard deviation of the predicted velocity trace, denoted $\sigma^{(w)}_{\widehat{v},k}$, with that of the ground–truth velocity trace, denoted $\sigma^{(w)}_{v,k}$. The directional consistency score is
$\mathrm{TCI}^{(w)}_{k}=1-\frac{\big|\sigma^{(w)}_{\widehat{v},k}-\sigma^{(w)}_{v,k}\big|}{\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}+\varepsilon},$
and is evaluated only when $\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}>\tau_{0}$, where $\tau_{0}=10^{-6}$ is a numerical tolerance that filters out angular directions with effectively no front motion, ensuring that TCI reflects meaningful temporal fluctuations rather than numerical artefacts. The final index aggregates all valid scores across windows and directions:
$\mathrm{TCI}=\frac{1}{W}\sum_{w=1}^{W}\frac{1}{\lvert\mathcal{K}_{w}\rvert}\sum_{k\in\mathcal{K}_{w}}\mathrm{TCI}^{(w)}_{k}\in[0,1],$
where $\mathcal{K}_{w}$ denotes the set of valid directions in window $w$; higher values indicate better preservation of intrinsic fluctuation dynamics.
To evaluate the organization of growth across angles, we examine the anisotropy of the radial velocity field. The normalized angular spread (NAS) is computed as the angular standard deviation divided by the angular mean of the velocity at each time point, where $\mathrm{NAS}_{t}$ refers to the value computed from the ground–truth velocity field and $\widehat{\mathrm{NAS}}_{t}$ refers to the value computed from the predicted one. We report their mean absolute deviation over time:
$\lvert\Delta\mathrm{NAS}\rvert=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{NAS}}_{t}-\mathrm{NAS}_{t}\big|.$
To further characterize directional patterning, we examine the angular Fourier spectrum of $v(\theta,t)$ and extract the normalized second–harmonic power,
$\mathrm{H}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[m]\big|^{2}},\qquad\widehat{\mathrm{H}}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[m]\big|^{2}},$
This quantity reflects the strength of the characteristic two–lobed angular mode. We define the Second–Harmonic Power Deviation as
$\lvert\Delta\mathrm{H}_{2}\rvert=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{H}}_{2,t}-\mathrm{H}_{2,t}\big|.$
4.7 Implementation of TexPol–Net
Framework and experimental setup. All models are trained using the Ultralytics YOLO framework, with the task configured for image segmentation. All experiments are conducted under identical training configurations and independently repeated to ensure fair comparability.
Training and optimization protocol. For training, the maximum number of epochs is set to 300, and an early stopping strategy is employed to mitigate overfitting once validation performance saturates. The batch size is set to 16, and all input images are resized to a fixed resolution of $640\times 640$, balancing computational efficiency and memory usage. Automatic mixed-precision (AMP) training is enabled throughout to improve training throughput and reduce memory consumption while maintaining numerical stability. The optimizer is set to stochastic gradient descent (SGD). The initial learning rate is set to $6\times 10^{-3}$ and is combined with a linear warm-up schedule over the first three epochs. After warm-up, the learning rate is gradually decayed to $1\%$ of its initial value. A weight decay of $5\times 10^{-4}$ and a momentum of $0.937$ are used to stabilize gradient updates and improve generalization. All experiments use a fixed random seed and enable deterministic training to ensure reproducibility.
4.8 Implementation of Morpher
Sequence construction and data preprocessing. Morpher operates exclusively on binary colony masks generated by a pretrained segmentation model and does not directly access raw image intensities. In our primary pipeline, TexPol–Net was used to generate these masks, while alternative segmentation models were used in comparative experiments. Each training sample therefore consisted of a temporally ordered sequence of segmentation masks representing colony occupancy and front geometry. Sequences were constructed by uniformly subsampling frames from the full temporal series using a fixed stride, yielding equal temporal spacing between adjacent frames, and the resulting sequence length $T$ was defined by this fixed-stride subsampling rule. Each sequence was partitioned into an observation segment and a prediction segment using observation–prediction ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$. Within each experiment, all masks were generated using the same segmentation model to ensure a consistent morphological representation across both segmentation and forecasting stages. All masks were resized to a spatial resolution of $640\times 640$, which was used consistently across all experiments. The dataset was split into training and validation partitions at the growth-sequence level, with no colony contributing sequences to both partitions, and all quantitative results are reported on the validation split.
Training and optimization protocol. All variants were optimized with AdamW using an initial learning rate of $5\times 10^{-5}$ and weight decay of $1\times 10^{-4}$. Training was conducted for 300 epochs with a batch size of 2. All input mask sequences were resized to a fixed spatial resolution of $640\times 640$. A linear warm-up was applied over the first 10% of optimization steps, followed by cosine annealing. The global gradient norm was clipped to 1.0. Mixed-precision training was enabled throughout via automatic mixed precision with gradient scaling to improve computational throughput while maintaining numerical stability. Validation was performed at every epoch, and model selection was based on the checkpoint achieving the highest validation mIoU. No early stopping was applied. All experiments were conducted with fixed random seeds and deterministic backend settings to ensure reproducibility. TensorFloat-32 acceleration was enabled for matrix multiplications on supported hardware, while cuDNN was configured in deterministic mode.
4.9 Running of existing methods
SAM and SAM2. SAM and SAM2 were fine-tuned for segmentation under a unified training protocol. For SAM, training was conducted using an adaptive optimizer with weight decay (initial learning rate $8\times 10^{-4}$; weight decay $1\times 10^{-4}$). A warm-up phase of $250$ optimization steps was applied at the beginning of training, followed by stepwise learning-rate decays at iteration steps 60,00060{,}000 and 86,66686{,}666, each with a decay factor of $1/10$. The model was built on the SAM ViT-B backbone, and a selective freezing strategy was adopted: the image encoder and prompt encoder were frozen, while only the mask decoder was updated during training. For SAM2, training was formulated as a binary segmentation task with RGB images as input and binary masks as supervision. Images were normalized to the range $[0,1]$ using ToTensor(), while masks were processed using nearest-neighbor interpolation to preserve label integrity. Masks were explicitly binarized to restrict pixel values to $\{0,1\}$. Optimization was performed using Adam with an initial learning rate of $1\times 10^{-4}$ and no additional learning-rate scheduling. A batch size of $4$ was used, and BCEWithLogitsLoss was applied directly to the raw logits.
MAU. MAU was run using the official implementation. The model employed four recurrent layers with hidden dimension 64, convolutional filters of size 5 with stride 1 and patch size 1, and no layer normalization. The spatiotemporal relation size was set to 2 and the temporal decay parameter to $\tau=5$. Scheduled sampling was enabled, with the sampling probability linearly decayed from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$. Training used the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a OneCycle learning rate scheduler.
MIM. MIM was run using the official implementation built on the PredRNN framework. The model employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 4, and no layer normalization. Scheduled sampling followed the same linear decay strategy as above, while reverse scheduled sampling was disabled. Training used the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a OneCycle learning rate scheduler; incomplete batches were dropped during training.
PredRNN and PredRNNv2. PredRNN-based models were run using the official implementations. Both models employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 2, and no layer normalization. Scheduled sampling was enabled with linear decay from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$. Training used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a OneCycle learning rate scheduler. PredRNNv2 additionally enabled reverse scheduled sampling with transition steps at 25,000 and 50,000 iterations and an exponential coefficient of 5,000, and incorporated a decoupling loss with weight $\beta=0.01$.
SimVP and SimVPv2. SimVP-based baselines were run using the official implementation. The spatial encoder–decoder employed a channel width of 64 with four convolutional blocks ($N_{S}=4$), while temporal modeling used a hidden dimension of 256 with eight temporal blocks ($N_{T}=8$). SimVP used TAU units for temporal prediction, whereas SimVPv2 replaced TAU with gSTA modules. Training used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a OneCycle learning rate scheduler. Model selection followed the validation loss criterion defined in the configuration.
4.10 Swarming Morphogenesis Evolution dataset
The Swarming Morphogenesis Evolution (SwarmEvo) dataset consists of high-resolution time-lapse recordings of Enterobacter sp. SM3 acquired at a fixed spatial resolution of $1250\times 1250$ pixels. After augmentation, the dataset comprises 1,971 annotated sequences used for training and evaluating segmentation models, as well as 276 long time series derived from continuous recordings sampled at 1-min intervals, which serve as the basis for temporal modeling and multi-scale temporal downsampling. Data were collected across multiple agar plates and independent imaging sessions, introducing natural variability in growth dynamics and colony morphology across distinct morphogenetic regimes. Segmentation masks used for model training and evaluation were obtained through a dedicated segmentation pipeline and subsequently curated to ensure consistent boundary delineation, while temporal sequences were generated by propagating these masks across time to support forecasting tasks. Details of cultivation conditions, imaging setup, and mask generation are provided in Section 4.1.
Segmentation-level augmentation For training the segmentation model, augmentations were applied independently to each image–annotation pair. Photometric perturbations included linear intensity rescaling with offset, gamma correction, additive Gaussian noise, and sparse impulse-like pixel corruption. Geometric transformations were sampled per image and applied consistently to the image and its polygon annotations, including random in-plane rotation, isotropic scaling, translation constrained by the instance extent, and random horizontal or vertical flipping. A random cutout was further used to simulate partial occlusion; polygon annotations were updated by geometrically clipping the visible region and retaining valid connected components. After each transformation, polygon validity was enforced by automatic closure and self-intersection repair, and invalid or degenerate shapes were discarded.
@@ Acknowledgements
S1 Performance comparison with state–of–the–art video prediction models under an 80%–20% observation–prediction split.
Table 1: Performance comparison with state–of–the–art video prediction models under an 80%–20% observation–prediction split. This table benchmarks Morpher against leading video prediction architectures, including MAU, MIM, PredRNN variants, and SimVP–based models. All methods are evaluated under identical input–output protocols for long–term forecasting of swarming colony expansion. Morpher achieves substantially higher region–level overlap (mIoU) and lower boundary error (HD95, ASSD), indicating improved accuracy in front propagation and boundary–level morphology.
Model mIoU (%) $\uparrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$
MAU [6] 84.67 22.73 14.68
MIM [46] 89.32 20.17 10.30
PredRNN [44] 84.60 22.75 14.81
PredRNNv2 [45] 84.14 23.24 15.04
SimVP [15]+TAU [40] 86.87 23.19 12.47
SimVP [15]+gSTA [39] 90.52 18.28 8.87
Morpher (Ours) 95.42 10.61 3.93
S2 Texture-Edge Attention and Polar–Context Attention modules
Swarming colony images exhibit complex morphological organization characterized by uncertain boundaries, irregular shapes, and radially propagating texture patterns. These inherent properties pose significant challenges for CNNs, whose reliance on local receptive fields restricts their capacity to capture long-range spatial dependencies and global geometric structures, particularly the near–concentric–ring radial expansion typical of swarming growth. To overcome these limitations and enhance both fine-grained texture extraction and high-level semantic representation, we developed two specialized attention modules, the Texture-Edge Attention (TEA) and Polar–Context Attention (PCA), as illustrated in Figure 6.
The TEA module, as shown in Figure 6LABEL:fig:TEA, is designed to address blurred boundaries and multi-scale, high-frequency texture variability. It combines three cooperative paths: a local branch that preserves intra-channel spatial details, a multi-dilated path that ensures scale-robust texture encoding, and an edge-sensitive path initialized with a discrete Laplacian kernel to enhance boundary awareness. Channel-wise and spatial gating mechanisms further refine the fused representation by emphasizing informative structures while maintaining computational efficiency.
@@
A squeeze–and–excitation (SE) gate $\mathbf{f}\!\in\!\mathbb{R}^{B\times C_{h}\times 1\times 1}$ is computed via global average pooling (GAP) followed by two $1\times 1$ convolutions with nonlinearity and sigmoid activation:
@@
Figure 6: Texture-Edge Attention (TEA) and Polar–Context Attention (PCA) modules. a, The TEA block enhances fine-scale texture fidelity and boundary sharpness through three cooperative branches: a local depthwise path for intra-channel spatial preservation, multi-dilated convolutions for scale-robust texture encoding, and an edge-sensitive Laplacian path that injects a high-pass prior. Channel and spatial gating further refine feature fusion, producing an edge-aware, redundancy-suppressed representation. b, The PCA block embeds a polar-aware geometric prior aligned with the radial growth of swarming colonies. Input features are first compressed and then processed by a local branch, a large-kernel Cartesian branch, and a polar-warped branch operating in $(\rho,\theta)$ coordinates. Depthwise dilated filters extract context along radial and angular axes, and subsequent channel– and spatial–attention gates yield a geometry-aligned output.
and concatenation of all paths yields
@@
where K=3K{=}3 captures short–, medium–, and long–range textures.
@@
While TEA focuses on boundary and texture fidelity, the PCA module illustrated in Figure 6LABEL:fig:PCA captures long-range dependencies and radial geometric organization inherent to swarming colonies. Conventional convolutional operators struggle with near–concentric–ring propagation whereas PCA embeds a polar-aware representation aligned with colony growth.
The module comprises three paths: a local branch for spatial detail, a large-kernel branch for contextual encoding, and a polar branch that transforms features into polar coordinates for radial modeling. The local branch follows the same $3\times 3$ depthwise–pointwise pattern described in Eq. 2, operating on $\mathbf{X}^{\prime}$:

### v2

@@ 5.3 Morpher for morphology forecasting
The temporal module is instantiated as one of four matched-capacity sequence models: vanilla RNN [16], GRU [13], LSTM [22], or Transformer encoder [59]. All variants share the same encoder–decoder backbone, the same latent dimensionality, and the same past-to-state formulation. The recurrent variants use three recurrent layers; the Transformer uses three encoder blocks.
Prediction is supervised at each forecast step by a temporally averaged objective,
$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}^{(t)},\quad\mathcal{L}^{(t)}=\lambda_{f}\mathcal{L}_{f}+\lambda_{i}\mathcal{L}_{i}+\lambda_{b}\mathcal{L}_{b}.$
Here $\mathcal{L}_{f}$, $\mathcal{L}_{i}$, and $\mathcal{L}_{b}$ denote focal, soft IoU, and boundary losses, respectively. All weights were fixed across experiments. Implementation details, sequence construction, and baseline-matched comparisons are provided in the Supplementary Information.
5.4 Training and implementation
For TexPol–Net, all images were resized to $640\times 640$. Training used the Ultralytics YOLO framework with SGD ($6\times 10^{-3}$ initial learning rate, three-epoch linear warm-up, decay to 1% of the initial rate, momentum =0.937=0.937, weight decay $5\times 10^{-4}$), batch size =16=16, mixed-precision training, and early stopping once validation performance saturated.
For Morpher, binary masks generated by a single segmentation model were uniformly subsampled with a fixed stride and partitioned into observation and prediction segments under ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$. All masks were resized to $640\times 640$. Training used AdamW ($5\times 10^{-5}$ initial learning rate, weight decay $10^{-4}$), batch size =2=2, 300 epochs, 10% warm-up followed by cosine annealing, gradient clipping at 1.0, and mixed precision. Model selection was based on the highest validation mIoU. No early stopping was applied. All experiments used fixed random seeds and deterministic backend settings; TensorFloat-32 acceleration was enabled where available.
All baseline models were retrained under matched preprocessing and comparable train/validation splits. Model-specific implementation details, including optimizer schedules and architecture-specific settings for YOLOv11/12, SAM/SAM2, MAU, MIM, PredRNN, PredRNNv2, SimVP+TAU, and SimVPv2+gSTA, are provided in the Supplementary Information.
5.5 Evaluation
Segmentation was evaluated by mAP50:95, image-wise IoU, and Dice coefficient. Forecasting was evaluated at two levels. Spatial fidelity was assessed by mIoU, HD, HD95, and ASSD, which quantify overlap and boundary localization. Dynamical fidelity was assessed by radial-velocity RMSE and the Temporal Consistency Index (TCI), which measure front propagation and temporal fluctuation preservation. Angular organization was quantified by $|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$, which measure deviations in anisotropic growth and second-harmonic angular structure.
In the main text, mAP50:95 serves as the primary segmentation metric, while HD95 and ASSD serve as the primary forecasting metrics because region overlap can remain high even when the contour drifts. Formal definitions of all metrics are provided in the Supplementary Information.
@@ Acknowledgements
Conflict of Interest
The authors declare no conflict of interest.
Data availability
All datasets used in this study are publicly available. The Swarming Morphogenesis Evolution (SwarmEvo) dataset, including both the segmentation and temporal prediction subsets, is hosted at https://huggingface.co/datasets/ShengyouDuan/SwarmEvo. The dataset provides polygon-based segmentation annotations and time-lapse sequences for bacterial swarming experiments, and is released to support reproducibility and further research in morphology-aware modeling.
Code availability
The complete implementation of the proposed framework, including TexPol–Net for morphology-aware segmentation and Morpher for autoregressive temporal forecasting, is publicly available at https://github.com/ShengyouDuan/From_shape_to_fate__making_bacterial_swarming_expansion_predictable. The repository contains all training and evaluation code required to reproduce the results reported in this work.
@@
S1 Texture–Edge Attention and Polar–Context Attention modules
Swarming colony images exhibit complex morphological organization characterized by uncertain boundaries, irregular shapes, and radially propagating texture patterns. These inherent properties pose significant challenges for CNNs, whose reliance on local receptive fields restricts their capacity to capture long-range spatial dependencies and global geometric structures, particularly the near-concentric-ring radial expansion typical of swarming growth. To overcome these limitations and enhance both fine-grained texture extraction and high-level semantic representation, we developed two specialized attention modules, the Texture–Edge Attention (TEA) and Polar–Context Attention (PCA), as illustrated in Figure 7.
The TEA module, as shown in Figure 7a, is designed to address blurred boundaries and multi-scale, high-frequency texture variability. It combines three cooperative paths: a local branch that preserves intra-channel spatial details, a multi-dilated path that ensures scale-robust texture encoding, and an edge-sensitive path initialized with a discrete Laplacian kernel to enhance boundary awareness. Channel-wise and spatial gating mechanisms further refine the fused representation by emphasizing informative structures while maintaining computational efficiency.
@@
A squeeze-and-excitation (SE) gate $\mathbf{f}\!\in\!\mathbb{R}^{B\times C_{h}\times 1\times 1}$ is computed via global average pooling (GAP) followed by two $1\times 1$ convolutions with nonlinearity and sigmoid activation:
@@
Figure 7: Texture–Edge Attention (TEA) and Polar–Context Attention (PCA) modules. a, The TEA block enhances fine-scale texture fidelity and boundary sharpness through three cooperative branches: a local depthwise path for intra-channel spatial preservation, multi-dilated convolutions for scale-robust texture encoding, and an edge-sensitive Laplacian path that injects a high-pass prior. Channel and spatial gating further refine feature fusion, producing an edge-aware, redundancy-suppressed representation. b, The PCA block embeds a polar-aware geometric prior aligned with the radial growth of swarming colonies. Input features are first compressed and then processed by a local branch, a large-kernel Cartesian branch, and a polar-warped branch operating in $(\rho,\theta)$ coordinates. Depthwise dilated filters extract context along radial and angular axes, and subsequent channel- and spatial-attention gates yield a geometry-aligned output.
The concatenation of all paths yields
@@
where K=3K{=}3 captures short-, medium-, and long-range textures.
@@
While TEA focuses on boundary and texture fidelity, the PCA module illustrated in Figure 7b captures long-range dependencies and radial geometric organization inherent to swarming colonies. Conventional convolutional operators struggle with near-concentric-ring propagation whereas PCA embeds a polar-aware representation aligned with colony growth.
The module comprises three paths: a local branch for spatial detail, a large-kernel branch for contextual encoding, and a polar branch that transforms features into polar coordinates for radial modeling. The local branch follows the same $3\times 3$ depthwise-pointwise pattern described in Eq. 2, operating on $\mathbf{X}^{\prime}$: